• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Fatiha

    Log entry time 00:01:53 on September 5,2009

    Entry number 288075

    keyword=Shift Summary

    Shift crew: F. Benmokhtar, K. Aniol and X. Zhan
    -----------

    IMPORTANT: Keep an eye on the ABU, the automated ABU values
    are going to the BANU values... Please fix it by hand until
    Bob gets the chance to have a look at it.

    16:00... Started shift with Bob doing the study of beam
    position/spot++ run

    16:18... Moved to Al dummy with a hole target, asked for 1uA.

    16:45... Moved to Al dummy, 1uA

    17:02... VDC HV turned off.

    17:08... Asked for 2 uA, with Al dummy, to check the spot size.

    17:16... Back to production, LH2, 100 uA

    Smooth production running for the rest of the shift.



    Run Summary:
    ------------
    Spectrometer DAQ runs:
    ----------------------
    24657/4466.... Spot test, 12C hole, 1uA --> well centered.
    24658/4467.... Spot test, Al dummy with a hole, ps factors are
    2 for both triggers. VDC HV = 0KV
    24659/4468.... Spot test, Al dummy with hole 1uA, VDC HV = 4KV
    500K event each.
    24660/4469.... Spot test, Al dummy, 1uA, VDC HV = 4KV
    24661 .... Spot++ check --> looks OK 5 by 4.5
    24662 .... Spot++, 100 uA, LH2 target

    Happex Runs:
    ------------
    12002... 100K
    12003... 100K
    12004... 100K
    12005... 100K
    12006... 100K
    12007... 100K
    12008...